OpenStudy (anonymous):

What is the slope of the line that passes through the pair of points (-5.2, 8.7) and (-3.2, 2.7)?

OpenStudy (gorv):

\[\frac{ 8.7-2.7 }{ -5.2+3.2 }=\frac{ 6 }{ -2 }=-3\]

OpenStudy (anonymous):

You make these look so easy. I am getting all messed up when I am dealing with numbers as negatives!